### v1

# From Introspection to Best Practices: Principled Analysis of Demonstrations in Multimodal In-Context Learning

Abstract
Motivated by in-context learning (ICL) capabilities of Large Language models (LLMs), multimodal LLMs with additional visual modality are also exhibited with similar ICL abilities when multiple image-text pairs are provided as demonstrations. However, relatively less work has been done to investigate the principles behind how and why multimodal ICL works. We conduct a systematic and principled evaluation of multimodal ICL for models of different scales on a broad spectrum of new yet critical tasks. Through perturbations over different modality information, we show that modalities matter differently across tasks in multimodal ICL. Considering such modality impact, we further utilize modality-driven demonstration strategies to boost ICL performance. We also identify that demonstration selection is closely related to the models’ ability to capture task inductive biases from multimodal ICL. Our principled analysis provides a comprehensive way of understanding the role of demonstrations in multimodal in-context learning, and sheds light on effectively improving multimodal ICL on a wide range of tasks even if those tasks are not seen in or even contradict pretraining data.
From Introspection to Best Practices: Principled Analysis of Demonstrations in Multimodal In-Context Learning
Nan Xu[Uncaptioned image]  Fei Wang[Uncaptioned image]  Sheng Zhang[Uncaptioned image]  Hoifung Poon[Uncaptioned image] Muhao Chen[Uncaptioned image] [Uncaptioned image]University of Southern California  [Uncaptioned image]Microsoft Research  [Uncaptioned image]University of California, Davis {nanx,fwang598}@usc.edu  {shezhan,hoifung}@microsoft.com  muhchen@ucdavis.edu
1 Introduction
(a) Questions and ground-truth answers from two of the investigated benchmarks: cross-style (left) and text-rich understanding (right).
(b) ICL performance against visual and textual perturbations.
(c) ICL performance given demonstrations selected by different modality-driven strategies.
Figure 1: (b) Modality matters differently on ICL across tasks: visual information matters little on Application but a lot on KIE, textual answers are more important to ICL on KIE than that to Application. (c) Demonstrations selected by text-driven strategy BERTScore benefits more on Application, while those selected by visual similarity (CLIP) bring higher accuracy on KIE.
Motivated by in-context learning (ICL) capabilities of Large Language Models (LLMs) for NLP tasks (Brown et al., 2020; Garg et al., 2022; Akyürek et al., 2022), multimodal LLMs with additional visual modality are also exhibited with similar ICL abilities when multiple image-text pairs are provided as demonstrations (Alayrac et al., 2022; Bai et al., 2023; Sun et al., 2023; McKinzie et al., 2024). In recent studies, the Retrieval-based In-Context Example Selection (RICES, Yang et al. (2022)) approach, which retrieves similar images in the support set by comparing their visual features with testing images, has become a default approach to select demonstrations for multimodal in-context learning (Alayrac et al., 2022; Sun et al., 2023; Yang et al., 2024).
However, relatively less work has been done to investigate the principles behind how and why multimodal ICL works, nor there has been enough justification for the necessity of selecting demonstrations according to visual modality and analyze its advantages over other modalities. Yang et al. (2024) only explored better in-context configurations for image captioning, while Chen et al. (2023) argued that multimodal ICL is predominantly driven by the textual information in the demonstrations. However, their observations are limited to image captioning (Young et al., 2014; Chen et al., 2015) and general-purpose visual question answering tasks (Goyal et al., 2017; Gurari et al., 2018; Marino et al., 2019; Sidorov et al., 2020), which leaves a comprehensive exploration for the strengths of ICL and its limitations (Zong et al., 2024) largely open for multimodal LLMs.
In this paper, we conduct a systematic and principled evaluation of multimodal ICL for models of different scales (ranging from OpenFlamingo 4B, Awadalla et al. (2023)) to IDEFICS1 80B, Laurençon et al. (2023)) on a broad spectrum of new yet critical tasks as shown in Figure 1(a). These tasks require different types of capabilities, including hallucination mitigation (Wang et al., 2023a), text-rich image understanding (Liu et al., 2023; Li et al., 2024), medical information comprehension (He et al., 2020; Pacheco et al., 2020; Liu et al., 2021), and cross-style transfer (Cai et al., 2023), etc. With diverse ICL capabilities examination, we show that the dependency of performance gain from ICL on demonstration modalities differs among tasks (Section 4). As demonstrated in Figure 1(b), perturbing visual information in demonstrations (e.g., removing or replacing with random, noised or permuted images) does not cause significant performance drop on ICL for cross-style and medical tasks, while resulting in decreased accuracy than that provided by correct demonstrations on tasks such as key information extraction from text-rich images. Sometimes it even leads to much worse performance than the zero-shot inference. On the other hand, textual perturbations (e.g., replacing the question/answer with random or one from other candidates in the same demonstration set) hurt ICL performance to different extents across tasks. For instance, perturbations on either questions or answers lead to greatly reduced accuracy on some tasks, while perturbations on answers results in extremely bad performance on others. These observations strongly suggest the necessity of understanding modality impact on ICL prior to collecting demonstrations for specific tasks.
We conduct further investigation on how to select effective demonstration to boost multimodal ICL performance (Section 5). As shown in Figure 1(c), we identify that providing demonstrations selected by textual similarity (e.g., text encoder of CLIP (Radford et al., 2021) or BERTScore (Zhang et al., 2019)) benefits ICL performance consistently across models and tasks. This is consistent with literature (Chen et al., 2023) and our prior observation that the textual modality plays an important role in ICL performance. For tasks observed with vital impact from visual modality on ICL performance, demonstrations selected by visual similarity (e.g., vision encoder of CLIP) elicit drastically improved ICL performance. Moreover, demonstration selection strategies that consider both visual and textual modalities, such as ALBEF (Li et al., 2021) with a multimodal encoder that explicitly models interactions between image and text features, present trade-off performance regardless of various modality importance to specific tasks.
Lastly, we illustrate that models may not always capture task inductive biases from multimodal ICL (Section 6). Concretely, we flip annotations of demonstrations to override semantic priors learned during pretraining (e.g., “Yes” to admit hallucinated objects in images and “No” to deny the presence actually existing objects in images). Small-scale models fail to comprehend or follow practices against prior knowledge provided by randomly sampled demonstrations. Surprisingly, models learn to follow inductive biases given demonstrations selected according to textual similarities, an emergent ability unlocked by scaling studied in literature (Wei et al., 2022b; Zhou et al., 2022). This is reasonable as flipped annotations mainly convey inductive biases through texts. Such capability of capturing inductive bias of demonstrations without scaling up models is more attractive than using semantic priors, since the model would be able to perform a wide range of tasks without further tuning, even if those tasks are not seen in or even contradict pretraining data (Wei et al., 2023).
In summary, our principled analysis provides a comprehensive way of understanding the role of demonstrations in multimodal in-context learning. We empirically show that (1) modalities matter differently in multimodal ICL across tasks (Section 4), (2) demonstration strategies considering modality impact are able to boost ICL performance (Section 5), (3) demonstration selection is closely related to models’ ability to capture task inductive biases from multimodal ICL (Section 6). Overall, our work aims to shed light on effectively improving multimodal ICL on a wide range of tasks even if those tasks are not seen in or even contradict pretraining data.
2 Related Work
Textual ICL
LLMs have been recognized as strong few-shot learners since their emergence (Brown et al., 2020). With ICL, LLMs are empowered to generalize to a wide range of tasks at inference even if those tasks are not seen in pretraining data (Garg et al., 2022; Akyürek et al., 2022). However, the performance of ICL is critically sensitive to the choices of demonstrations (Rubin et al., 2022; Wang et al., 2023b; Gupta et al., 2023), the order (Lu et al., 2022; Wu et al., 2023) and format of prompts (Zhao et al., 2021; Min et al., 2021).
To understand why ICL works, Xie et al. (2021) explained from the theoretical perspective that Transformers acquire the ICL ability when they are trained to infer latent concepts during pretraining. Min et al. (2022) empirically showed that the performance gain of ICL over zero-shot inference is mainly driven by the label space, distribution of input text, output labels, and overall format of the sequence, while the represented mapping from inputs to the outputs in demonstrations matters little. However, some recent work (Zhou et al., 2022; Wei et al., 2023) suggested that when scaling up to some extent, larger models (e.g. PaLM-540B (Chowdhery et al., 2023) and Codex (Chen et al., 2021)) can actually learn input-output mappings, which allows them to perform a variety of challenging tasks even if they contradict pretraining data.
Considering the additional visual information in multimodal ICL, we study the importance of different modalities and guide demonstration selection for better ICL performance accordingly.
Multimodal ICL
After pretraining on interleaved image-text data or fine-tuning on multi-turn conversations, multimodal LLMs have exhibited ICL abilities in tasks such as image captioning and general-purpose visual question answering (Alayrac et al., 2022; Bai et al., 2023; Sun et al., 2023; McKinzie et al., 2024). Considering these studies may not sufficiently reveal strengths and weaknesses of ICL, Zong et al. (2024) recently introduced VL-ICL Bench which encompasses a broad spectrum of tasks for multimodal ICL evaluation. However, there is not much work that conducts principled analysis on emergent ICL capabilities and provides insightful suggestions for future ICL practices. Yang et al. (2024) only explored better in-context configurations for image captioning.
One work that is closely connected to ours is Chen et al. (2023). Chen et al. (2023) argued that multimodal ICL is predominantly driven by the textual information in the demonstrations and proposed Mixed Modality In-Context Example Selection (MMICES), which first pre-filters samples based on visual feature similarity and then selects most similar ones based on textual similarity. However, their observations are limited to image captioning and general-purpose visual question answering tasks, which leaves a comprehensive exploration for the strengths of ICL and its limitations (Zong et al., 2024) largely open for multimodal LLMs. We conduct more comprehensive study on the impact of modality to ICL and find that modalities matter differently across tasks. Furthermore, we investigate how models of different scales capture task inductive biases from multimodal ICL.
3 Experimental Setup
Capabilities Tested Dataset #Train #Test Metric References
Captioning Image COCO 2,815,816 500 CIDEr Chen et al. (2015)
Flickr30K 29,000 500 CIDEr Young et al. (2014)
General visual perception and textual understanding OKVQA 9,009 500 Accuracy Marino et al. (2019)
VQAv2 443,757 500 Accuracy Goyal et al. (2017)
TextVQA 34,602 500 Accuracy Sidorov et al. (2020)
VizWiz 20,523 500 Accuracy Gurari et al. (2018)
In-context Learning VL-ICL 9,960 1,120 Accuracy Zong et al. (2024)
Mathematical Reasoning MATH-Vision 2,540 500 Accuracy Wang et al. (2024)
Hallucination AMBER Existence 8,763 500 Accuracy Wang et al. (2023a)
AMBER Attribute 7,124 500 Accuracy
AMBER Relation 1,163 500 Accuracy
Text-rich Visual Comprehension OCRBench 53,991 900 Accuracy Liu et al. (2023)
SEED-Bench-2-Plus 1,174 1,103 Accuracy Li et al. (2024)
Medical Path-VQA 19,755 500 Token F1 He et al. (2020)
Slake-VQA 9,835 500 Token F1 Liu et al. (2021)
PAD-UFES-20 994 500 Accuracy Pacheco et al. (2020)
Multiple Images Seed-Bench-2 3,751 2,260 Accuracy Li et al. (2024)
Cross-style BenchLMM Artistic 100 400 Accuracy Cai et al. (2023)
BenchLMM Sensor 300 400 Accuracy
BenchLMM Application 367 400 Accuracy
Table 1: Evaluation benchmark statistics. We adopt the default train and test split as the demonstration candidates and testing examples if the testing annotations are provided, otherwise the validation split is used instead. We randomly sample at most $500$ instances for testing. The three datasets marked by  are composed of multiple subsets and we consider average performance for analysis, leaving detailed results in Appendix.
Multimodal LLMs Visual Encoders LLMs #Params
openai CLIP ViT-L/14 togethercomputer/RedPajama-INCITE-Base-3B-v1 4B
OpenFlamingo-4B https://huggingface.co/openflamingo/OpenFlamingo-4B-vitl-rpj3b
openai CLIP ViT-L/14 anas-awadalla/mpt-7b 9B
OpenFlamingo-9B https://huggingface.co/openflamingo/OpenFlamingo-9B-vitl-mpt7b
laion/CLIP-ViT-H-14-laion2B-s32B-b79K huggyllama/llama-7b 9B
IDEFICS1-9B https://huggingface.co/HuggingFaceM4/idefics-9b
laion/CLIP-ViT-H-14-laion2B-s32B-b79K huggyllama/llama-65b 80B
IDEFICS1-80B https://huggingface.co/huggyllama/llama-65b
IDEFICS2-8B google/siglip-so400m-patch14-384 mistralai/Mistral-7B-v0.1 8B
https://huggingface.co/HuggingFaceM4/idefics2-8b-base
EVA-CLIP LLaMA 14B
Emu1 https://huggingface.co/BAAI/Emu/blob/main/Emu-pretrain.pt
Table 2: Information of tested multimodal LLMs, their visual encoder, text models, number of parameters and the download links on Hugging face.
In this section, we describe the experimental setup used in our analysis (Section 4-Section 6). We list evaluation benchmarks and corresponding metrics in Table 1, as well as studied model information in Table 2.
Evaluation Benchmarks
After pretraining multimodal LLMs on interleaved image-text data or fine-tuning on multi-turn conversations, existing work (Alayrac et al., 2022; Bai et al., 2023; Sun et al., 2023; McKinzie et al., 2024) mainly focuses on evaluating their in-context learning abilities on image captioning such as COCO (Chen et al., 2015) and Flickr30K (Young et al., 2014), as well as general-purpose visual question answering tasks such as OKVQA (Marino et al., 2019), VQAv2 (Goyal et al., 2017), TextVQA (Sidorov et al., 2020) and VizWiz (Gurari et al., 2018). Besides these classic vision-language tasks, we also consider one recently released benchmark, namely VL-ICL Bench (Zong et al., 2024), which encompasses a broad spectrum of challenging new tasks to investigate strengths and limitations of in-context learning capabilities.
Benefits of utilizing demonstrations as contexts for more critical and practical applications, though imperfect zero-shot performance is observed from state-of-the-art models, are not yet explored. Therefore, we further study in-context learning capabilities of multimodal LLMs on the following tasks. 1) Math Reasoning: MATH-Vision (Wang et al., 2024) is a large math reasoning benchmark that collects questions from real math competitions and tests the general visual perception and mathematical reasoning abilities; 2) Hallucination: AMBER Wang et al. (2023a) provides a discriminative way to evaluate various types of hallucination including existence, attribute and relation; 3) Text-rich Tasks: both OCRBench (Liu et al., 2023) and SEED-Bench-2-Plus (Li et al., 2024) assess text-rich visual comprehension of models, while the former focus on Optical Character Recognition (OCR) capabilities and the latter covers text-rich scenarios in the real world such as Charts, Maps, and Webs; 4) Medical Tasks: three datasets consider different medical modalities, i.e., Path-VQA (He et al., 2020) for pathology, Slake-VQA (Liu et al., 2021) for radiology and PAD-UFES-20 (Pacheco et al., 2020) for skin lesion images. 5) Multi-image Tasks: Seed-Bench-2 Li et al. (2024) evaluates the ability to comprehend multimodal inputs containing multiple images. 6) Cross-style Transfer: BenchLMM (Cai et al., 2023) assesses the robustness of models against three different styles including artistic image, imaging sensor, and application styles.
Multimodal LLMs
We evaluate pretrained multimodal LLMs without further instruction tuning, so that factors, such as seeing similar data or acquiring tested capabilities from the instruction dataset rather than through in-context learning, could be fairly reduced. Specifically, we consider the following pretrained models that scale from 4B to 80B and have previously demonstrated in-context learning abilities through limited analysis: OpenFlamingo (Awadalla et al., 2023) of two sizes (4B and 9B), IDEFICS of two scales from different versions (9B and 80B from the 1st version (Laurençon et al., 2023) and 8B from the 2nd version Laurençon et al. (2024)), together with the 14B Emu1 (Sun et al., 2023).
Moreover, we evaluate the proprietary model, GPT-4o (OpenAI, 2024), to exhibit challenge levels of evaluated tasks on the one hand, and compare in-context learning capabilities between pretrained and instruction-tuned models on the other hand.
Evaluation Metrics
For image captioning, we report CIDEr (Vedantam et al., 2015) scores. For general-purpose VQA tasks, we adopt the common VQA evaluation metric (Antol et al., 2015), where $10$ annotations are provided and the model prediction is deemed $100\%$ accurate if at least three annotators provided that exact answer. To evaluate performance on two medical VQA task-slake-VQA and Path-VQA, we use the token-level F1 score following Tu et al. (2024). We follow the evaluation practices in BenchLMM where ChatGPT is employed to gauge the proximity of answers predicted by the LMMs to ground-truth answers. For remaining datasets, we utilize their original evaluation strategy–soft string matching, to eliminate the impact of answer formats.
Implementation Details
We prompt multimodal LLMs with an instruction “Describe the image:” for caption generation, while employing open-ended answer generation for other tasks with a prompt in the form of “Question: the <question> Answer:”, without any constraint on model’s output space.  [^1] For short answer generation, we modify the prompt slightly to “Question: <question> Short answer:”. We adopt the default decoding strategy and configurations (e.g., beam search with 5 as the number of beams for Emu1) suggested by each model vendor respectively. In contrast to the zero-shot setting, we consider 4- and 8-shot for in-context learning analysis  [^2] Considering limited amounts of images per example used for pretraining, we evaluate 1- and 2-shot performance on tasks from SEED-Bench-2 where each example contains at least 8 images., where the demonstrations are randomly sampled from candidates for each testing example unless otherwise stated.  [^3] For each testing example, the demonstrations are randomly sampled from the train set while shared among all studied models.
4 Modalities Matter Differently in Multimodal ICL
(a) Visual Modality
(b) Textual Modality
Figure 2: Multimodal ICL performance of IDEFICS1-80b reacts differently across tasks of different difficulty levels against perturbations on visual (top) and textual (bottom) information. For easy (i.e., BenchLMM Sensor and Application) and two moderate (i.e., Path-VQA and Slake-VQA) tasks, performance after various visual perturbations is very close to that given original correct demonstrations, while drops obviously when either textual question or answer is perturbed. For the moderate PAD-UFES-20, neither of the two modalities matter too much. For the hard KIE, we observe degraded performance when the image is removed or replaced. Similar observations from other 5 models can be found from Figure 13 to Figure 17.
As shown in Figure 5 and Footnote 7, pretrained models and GPT-4o generally achieve better performance given demonstrations as context in existing ICL tasks. As demonstrated in Figure 7, on more complex and reasoning-focused tasks, pretrained models generally benefit more from demonstrations while the performance of GPT-4o is barely influenced.
In this section, we examine which modality of the demonstrations takes more effect in multimodal in-context learning. For a comprehensive evaluation, we focus on three tasks of different difficulty levels: easy cross-style tasks (i.e., BenchLMM Sensor and Application in Figure 7), moderate medical tasks (i.e., Path-VQA, Slake-VQA and PAD-UFES-20 in Figure 7), and hard text-rich key information extraction task (i.e., KIE from OCRBench in Figure 10). We visualize $4$-shot performance of IDEFICS-80b within this section while leaving results of other models in Appendix (from Figure 13 to Figure 17.). We identify that the dependency of performance gain from ICL on demonstration modalities differs among tasks.
33footnotetext: In zero-shot setting, GPT-4o achieves extremely poor performance on general-purpose VQA datasets such as OKVQA, VQAv2, TextVQA and VizWiz. We find that GPT-4o tends to provide long answers even after we give the instruction “Always provide short answers.” as the system message. This results in low scores when comparing against short annotations.
4.1 Impact of Visual Modality
In recent studies, the Retrieval-based In-Context Example Selection (RICES (Yang et al., 2022)) approach, which retrieves similar images in the support set by comparing their visual features with testing images, has become a default approach to select demonstrations for multimodal in-context learning (Alayrac et al., 2022; Sun et al., 2023; Yang et al., 2024). However, the necessity of selecting demonstrations according to visual modality and its advantages over other modalities are not yet explored.
By fixing the textual modality (i.e., question and answer pairs) of demonstrations, we experiment with demonstrations containing different perturbations of visual modality: 1) no images where only textual question and answer pairs are provided; 2) zero/one images that all zero (black)/255(white) pixel values are used instead; 3) noised images that apply Gaussian noises to the original images; 4) random images sampled from the train set; 5) permuted images reorganize the order of demonstration images so that visual and textual modalities are misaligned.
Results
We compare ICL performance of IDEFICS1-80B  [^4] We show performance of IDEFICS1-80B in  Figure 2 on all tasks except KIE, which is too challenge for IDEFICS1-80B to handle in both zero- and few-shot settings (at most 3 of 200 testing examples are answered correctly). Only IDEFICS2-8b can solve considerable amounts of cases (30 out of 200 in 4-shot setting), hence we perturb modality information on IDEFICS2-8b instead. before and after visual perturbations in Figure 2 and other models from Figure 13(a) to Figure 17(a). For easy cross-style and moderate medical tasks, we find that perturbing visual information in demonstrations does not cause significant performance drop on ICL, which is consistent with observations from prior work (Chen et al., 2023). However, for the hard KIE task, visual perturbations that remove or change content of images result in decreased accuracy than that provided with correct demonstrations, sometimes much worse performance than the zero-shot inference. This indicates that visual information plays an important role in improving ICL performance over zero-shot one, which is reasonable since this dataset requires extracting key-value pairs in the image (Liu et al., 2023). Meanwhile, the performance after applying Gaussian noises to images is very close to performance with correct images, which implies that multimodal LLMs are agnostic to image noises and able to extract key visual information for question answering.
4.2 Impact of Textual Modality
Previous studies have identified excessive dependence of multimodal LLMs on the language model’s linguistic priors (Han et al., 2022; Li et al., 2023). Accordingly, the role of textual modality for multimodal ICL should be similarly important. Therefore, we keep the visual modality of demonstrations while perform the following perturbations upon textual question and answer pairs: 1) no questions/answers remove the question/answer component directly; 2) random questions/answers employs questions/answers sampled from the train set instead; 3) permuted questions/answers exchange question or answer component of demonstration examples while keep the other two components unchanged.
Results
In Figure 2(b), we visualize ICL performance in response to perturbations upon questions or answers of demonstrations independently. We find that textual perturbations hurt ICL performance to different extents. On tasks such as BenchLMM Sensor, Slake-VQA and KIE, perturbations on either questions or answers lead to greatly reduced accuracy even below zero-shot inference. By replacing correct answers from demonstrations with random ones or those misaligned with image-question pairs, we observe extremely bad performance on Slake-VQA and KIE. On other tasks, questions and answers are almost equally important to exhibited ICL performance.
5 How to Select Effective Demonstrations for Multimodal ICL
Figure 3: Influence of modality-driven demonstration selection strategies on ICL performance of IDEFICS1-80B. Text-driven demonstration (e.g.,   textual CLIP,   BERT, and   BERTScore) selection strategies always bring performance improvement over zero-shot inference and random strategy. Strategies considering visual modality (e.g.,   visual CLIP,   MMICES, and   ALBEF) enhance performance significantly on KIE, where visual modality proves to be critical for ICL performance as illustrated in Figure 2(a). We visualize similar observations of other five models from Figure 13 to Figure 17.
Motivated by variational roles of different modalities across different tasks, we further explore influence of modality-driven demonstration selection strategies on ICL performance in this section.
Vision-driven Demonstration Selection
To retrieve demonstrations containing images similar to those in testing examples, we follow prior studies (Alayrac et al., 2022; Sun et al., 2023; Yang et al., 2024) by adopting the RICES strategy (Yang et al., 2022), which compares visual similarity according to features extracted from the pretrained visual encoder of CLIP (Radford et al., 2021).
Text-driven Demonstration Selection
For fair comparison with RICES, we employ the textual encoder of CLIP as well for selecting demonstrations with similar textual features to testing examples. We also adopt the BERTScore (Zhang et al., 2019) metric  [^5] We adopt the DeBERTa large model fine-tuned with MNLI task, which is accessible at https://huggingface.co/microsoft/deberta-large-mnli., which considers token-level similarity between candidate and reference sentences and shows strong correlation with human judgements on multiple common benchmarks.
Dual-modality driven Demonstration Selection
We first consider Mixed Modality In-Context Example Selection (MMICES) proposed by Chen et al. (2023), which first pre-filters $K$ samples ($K$=32) based on visual feature similarity and then selects most similar ones based on textual similarity. To represent vision-language features, we utilize ALBEF (Li et al., 2021), a multimodal encoder that explicitly models the interactions between image and text features and achieves state-of-the-art performance on image-text retrieval tasks. Since its multimodal encoder is built upon an image encoder (i.e., visual transformer ViT-B/16) and a text encoder (i.e., $\text{BERT}_{\text{base}}$), we also select demonstrations according to the embedding of the [CLS] token from $\text{BERT}_{\text{base}}$ as another textual-driven approach for contrast. For fair comparison, the vision-driven CLIP approach, the visual feature extractor of MMICES, and the visual encoder of ALBEF share the same visual transformer (i.e., ViT-B/16).
We focus on demonstration selection in this paper. Considering the sensitivity of LLMs to the ordering in the prompt (Lu et al., 2022; Wu et al., 2023), we follow prior work (Alayrac et al., 2022; Gupta et al., 2023) with demonstrations ordered by an increasing order of similarity, such that the most similar demonstration appears right before the testing example.
Results
We illustrate influence of demonstration selection strategies on ICL performance in Figure 3. Providing demonstrations selected by textual similarity benefits ICL performance consistently across models and tasks. This is consistent with literature (Chen et al., 2023) and our observations in Section 4.2 that the textual modality plays an important role in ICL performance. In general, the larger text embedding model–BERTScore (124M parameters) leads to better ICL performance compared with smaller models like textual CLIP (63M parameters) and BERT (124M parameters).
As analyzed in Section 4.1, visual information of demonstrations is of vital importance to ICL performance for the task KIE that requires key-value pair extraction from images. Accordingly, we witness drastically improved ICL performance when demonstrations containing more similar images to testing images are provided by visual CLIP to multimodal LLMs.
Strategies that consider dual modalities for demonstration selection (e.g., MMICES and ALBEF) are similarly more advantageous compared with text-driven methods on KIE. We also find that they achieve trade-off performance regardless of various modality importance to specific tasks. Meanwhile, ALBEF that explicitly models the interactions between image and text features obtains better ICL performance than MMICES, which is constrained by the vision-driven pre-filter process.
6 Models May Not Always Capture Task Inductive Biases from Multimodal ICL
Figure 4: The ability to capture inductive biases that contradict semantic priors when presented with flipped in-context exemplar annotations of AMBER Attribution emerges when demonstrations are selected according to textual modality (i.e.,   textual CLIP,   BERT and   BERTScore). Ground truth annotations for testing examples are not flipped, so if a model learns to follow flipped labels in demonstrations, its accuracy should be below $50\%$. Given random demonstrations   or those selected considering visual modality  , models cannot flip predictions to follow flipped annotations, while models can do so provided with demonstrations selected by text-driven strategies (performance decreases to well below $50\%$). We show similar observations on AMBER Existence and Relation in Figure 18.
Prior work on NLP tasks shows that small language models like GPT-J-6B (Wang and Komatsuzaki, 2021), PaLM-8B (Chowdhery et al., 2023) and GPT3 curie-6.7B (Gao et al., 2021) rely primarily on semantic priors from pretraining (Min et al., 2022), while large models such as PaLM-540B, InstructGPT (Ouyang et al., 2022) and Codex (Chen et al., 2021) can capture and follow inductive biases from in-context exemplars even when they contradict strong semantic priors that larger models may hold (Wei et al., 2023). However, it is unknown whether capturing inductive biases is still an emergent ability of model scale for multimodal ICL. In this section, we experiment with flipped labels on the $8$-shot hallucination benchmark AMBER–“Yes” is provided as demonstration annotation if the existence/attribute/relation description in the question is WRONG according to the image, “No” otherwise. We investigate both random and distinct modality-driven demonstration selection strategies to analyze the relation of capturing inductive biases from ICL to model scales and demonstration quality.
Results
We show the abilities of different models for capturing inductive biases from demonstrations in Figure 4. We flip annotations of demonstrations while keeping the ground-true answers of testing examples unflipped, hence the lower the accuracy, the stronger capabilities of multimodal LLMs to capture inductive biases and further override semantic priors learned during pretraining. When provided with demonstrations randomly sampled or selected according to similarities of visual features (i.e., visual CLIP), all evaluated models fail to comprehend or follow practices against prior knowledge. This is consistent with existing studies showing that small language models ignore flipped labels presented in-context and thus rely primarily on semantic priors from pretraining (Wei et al., 2023). Surprisingly, all studied small-scale models tend to follow inductive biases from demonstrations with accuracy well below $50\%$ when we switch demonstrations to those selected according to textual similarities (e.g., textual CLIP, BERT, BERTScore). We suspect that flipped annotations mainly convey inductive biases through texts, which makes text-driven selection strategies effective in guiding the behavior of small models to override semantic priors.
Notably, GPT-4o always follows the strong semantic priors and provide factual responses even when the demonstration annotations are flipped, which is quite opposite to emergent ability unlocked of model scale discovered in the literature (Wei et al., 2022a, 2023). However, GPT-4o’s failure to provide flipped answers following demonstrations does not indicate such large model is unable to capture those inductive biases. We speculate that GPT-4o may be able to perceive provided biases that are against semantic priors, but reject to give non-factual responses due to its built-in safety mechanisms across modalities (OpenAI, 2024).
7 Conclusion
We conduct a systematic and principled evaluation of multimodal ICL for models of different scales on a broad spectrum of new yet critical tasks. We find that modalities matter differently in multimodal ICL across tasks. Hence we utilize modality-driven demonstration strategies to boost ICL performance. Moreover, we find that demonstrations selected according to textual similarity help models capture inductive biases from multimodal ICL.
Limitations
We conduct a systematic and principled evaluation of multimodal ICL for pretrained models of different scales on a broad spectrum of new yet critical tasks. One limitation of our study is lack of discussion over instruction-tuned models, which may present differently than pretrained ones.
Ethics Statement
This paper presents comprehensive study of multimodal ICL on multiple existing benchmarks that have gone through ethical reviews in prior works. Therefore, we believe our work does not pose additional ethical issues.
Appendix A Appendix
Figure 5: Evaluation on existing ICL tasks including image captioning (COCO and Flickr30K), general-purpose VQA (OKVQA, VQAv2, TextVQA and VizWiz) and recently released VL-ICL benchmark. We observe ICL abilities in general, with different levels across tasks, models and demonstration amounts: Emu1 benefits less from provided demonstrations compared with others, while two OpenFlamingo models exhibit worse performance when 8 demonstrations are provided. Refer to Footnote 7 for GPT-4o performance and  Figure 8 for detailed results on each subset of the VL-ICL benchmark.
Figure 6: Performance of GPT-4o on existing ICL benchmarks. GPT-4o obtains much better performance when more demonstrations are given as the context.77footnotemark: 7
(a) Pretrained Models
(b) GPT-4o
Figure 7: More comprehensive ICL capability evaluation of pretrained models (top 3 row) and GPT-4o (bottom row) on recently proposed benchmarks. Pretrained models exhibit ICL abilities across different tasks, while GPT-4o achieves much higher zero-shot performance but benefits merely from provided demonstrations.
Figure 8: Detailed evaluation on VL-ICL Bench for multimodal in-context learning.
Figure 9: Detailed performance of GPT-4o on VL-ICL Bench for multimodal in-context learning.
Figure 10: Detailed evaluation on OCRBench (top) and SEED-Bench-2-Plus (bottom) for accessing Optical Character Recognition (OCR) capabilities.
Figure 11: Detailed evaluation on SEED-Bench-2 for the ability to comprehend multiple images and texts.
Figure 12: Detailed evaluation of GPT-4o on OCRBench (top), SEED-Bench-2-Plus (middle) and SEED-Bench-2 (bottom).
(a) Visual Modality
(b) Textual Modality
(c) Textual Modality
Figure 13: ICL performance of OpenFlamingo2-4B against perturbations over visual (top) and textual (bottom) modalities and different demonstration selection strategies (bottom).
(a) Visual Modality
(b) Textual Modality
(c) Textual Modality
Figure 14: ICL performance of IDEFICS2-8b against perturbations over visual (top) and textual (bottom) modalities and different demonstration selection strategies (bottom).
(a) Visual Modality
(b) Textual Modality
(c) Textual Modality
Figure 15: ICL performance of OpenFlamingo-9B against perturbations over visual (top) and textual (bottom) modalities and different demonstration selection strategies (bottom).
(a) Visual Modality
(b) Textual Modality
(c) Textual Modality
Figure 16: ICL performance of IDEFICS1-9b against perturbations over visual (top) and textual (bottom) modalities and different demonstration selection strategies (bottom).
(a) Visual Modality
(b) Textual Modality
(c) Textual Modality
Figure 17: ICL performance of Emu1-14B against perturbations over visual (top) and textual (bottom) modalities and different demonstration selection strategies (bottom).
(a) AMBER Existence
(b) AMBER Relation
Figure 18: The abilities to capture inductive biases with flipped in-context annotations on AMBER Existence and Relation dataset.